A thin disc of radius R and mass M has charge q uniformly distributed on it. It rotates with angular velocity w. The ratio of magnetic moment and angular momentum for the disc is(1) q/2M (2) R/2M (3) q^2/2M (4) 2M/q

Grade:12th pass

3 Answers

Arun
6 years ago
Dear student
\gamma = \mu /L
 = i * \pi * r²/m r² w = (q *w* \pi /2\pi) /mw = q/2m
Hence option 1 is correct.
Regards
Arun (askIITians forum expert)
Satyam
5 years ago
M= q×R^2×w/4      ........1
L= mR^2/2      ...........2
 By dividing equation 1 by 2 we will gell,
M/L= q/2m
[M=magnetic moment, w= angular velocity, m= mass, R= rafius, L=angular mementum]
Rishi Sharma
4 years ago
Dear Student,
Please find below the solution to your problem.

M= q×R^2×w/4 ........1
L= mR^2/2 ...........2
By dividing equation 1 by 2 we will gell,
M/L= q/2m
[M=magnetic moment, w= angular velocity, m= mass, R= rafius, L=angular mementum]

Thanks and Regards